1. The sum of all coefficients in the reaction H2O + P2O5 H3PO4
is:
A. 8 C. 4
B. 6 D. 10
1
Expert's answer
2020-05-16T13:42:39-0400

"3H_2O+P_2O_5=2H_3PO_4"

The sum=3+1+2=6

Answer: B.6